OpenStudy (anonymous):

in the system of x+2y=9 and x-y=6. what would be substituted for x and the second equation

OpenStudy (nubeer):

rear-rage your first equation x = 9-2y right?

OpenStudy (nubeer):

*rearrange

OpenStudy (anonymous):

it was wrong

OpenStudy (nubeer):

what is wrong about it?

OpenStudy (jiteshmeghwal9):

first find the value of 'x' from one of the equation \(x+2y=9\) \(x=9-2y\) now substitute he value in next equation & solve for 'y' :)

OpenStudy (jiteshmeghwal9):

after getting the value of 'y' put the value in the equation \(x=9-2y\) to gt the value of x :)

OpenStudy (anonymous):

you have to solve x in the second equations

OpenStudy (jiteshmeghwal9):

(9-2y)-y=6

OpenStudy (jiteshmeghwal9):

solve this equation for y :)

OpenStudy (anonymous):

no your soving for x and the second equation

OpenStudy (jiteshmeghwal9):

plz i m helping u. do it as i say :)

OpenStudy (jiteshmeghwal9):

solve for y:- (9-2y)-y=6 9-2y-y=6 9-3y=6 -3y=6-9 -3y=-3 y=1

OpenStudy (jiteshmeghwal9):

substituting the value of y in the equation x=9-2y x=9-2(1)=7

OpenStudy (jiteshmeghwal9):

so, x=7 & y=1 :)